OpenStudy (anonymous):

Use the Pythagorean Theorem to determine if the three numbers could be the measures of the sides of a right triangle. Assume that the third measure is the hypotenuse. yes no @Jamierox4ev3r I have no idea how to work it with the square on the end? 0.0

OpenStudy (anonymous):

if a^2 + b^2 = c ^2 ??

HanAkoSolo (jamierox4ev3r):

what are the measures anyhow @GreenBeanDemphsey lol!

OpenStudy (anonymous):

sorry i forgot to put them, 4, 7, and square 65

OpenStudy (anonymous):

square 65?

HanAkoSolo (jamierox4ev3r):

do you mean \(\sqrt{65}\)?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Square the entire equation then \[4 ^{2} + 7 ^{2} =? 65\]

HanAkoSolo (jamierox4ev3r):

because 16+49= what? ;) Hint: \(4^2=16~ and~ 7^2=49\)

OpenStudy (anonymous):

16 + 49 | 65

OpenStudy (anonymous):

65 = 65

OpenStudy (anonymous):

65 i know that

OpenStudy (anonymous):

i know how to add them and everything but how do i do it with the root

OpenStudy (anonymous):

So yes, the answer is yes, it's a right triangle

HanAkoSolo (jamierox4ev3r):

right so there is your answer! because that just affirms the equation and also affirms that this does indeed equal a right triangle :)

OpenStudy (anonymous):

When you square a square root, it becomes normal again

OpenStudy (anonymous):

okay, to the answer would be true

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

thank you everyone, lol!